OpenStudy (anonymous):

solve for θ. sin^2 θ= cos^2 θ

OpenStudy (anonymous):

so \(\sin(\theta)=\cos(\theta)\) or \(\sin(\theta)=-\cos(\theta)\) try\(\frac{\pi}{4},\frac{3\pi}{4}...\)

OpenStudy (anonymous):

I'm a bit confused isn't it 1/2(1-cos ^2θ) = 1/2(1+ cos^ θ)

OpenStudy (anonymous):

\[a^2=b^2\implies a=b\text{ or }a=-b\]

OpenStudy (anonymous):

\[\sin^2\theta+\cos ^{2}\theta = 1\]

OpenStudy (anonymous):

maybe i am reading it wrong thought it was \[\sin^2(\theta)=\cos^2(\theta)\]

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok then my answer above should be correct

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{\pi}{4}\] they are the same, also at \(\frac{5\pi}{4}\) at \(\frac{3\pi}{4}\) one is the negative of the other, also at \(\frac{7\pi}{4}\)

OpenStudy (noelgreco):

Divide both sides by cos^2(x) tan^2(x)=1 tan(x) = +1, -1

OpenStudy (anonymous):

I got + and - pi/4

OpenStudy (anonymous):

all of these answers are probably correct, just a lot of different ways to solve this problem

OpenStudy (anonymous):

there are an infinite number of answers, but in the interval \([0,2\pi)\) there are 4

OpenStudy (anonymous):

yeah, i see that now, good job